# Deep Learning for Continuous-Time Stochastic Control with Jumps

### v2

@@ 5 Numerical experiments
In our numerical experiments, we choose the DGM architecture of Sirignano & Spiliopoulos (2018) for both the value and optimal control networks, as it has been shown to empirically improve the PINN performance. Details on the network design and hyperparameters are given in Appendix B.
@@ 5.1 Linear-quadratic regulator with jumps
$\inf_{\alpha}\mathbb{E}\left[\int_{0}^{T}c_{1}\,\|\alpha_{s}\|^{2}_{2}\,ds+c_{2}\|X^{\alpha}_{T}\|^{2}_{2}\right],\vskip-0.28453pt$ (25)
where the infimum is over $d$-dimensional predictable
@@
processes $(\alpha_{t})_{0\leq t\leq T}$ and $X^{\alpha}$ is a $d$-dimensional process with controlled dynamics
$dX^{\alpha}_{t}=\alpha_{t}\,dt+\Sigma\,dW_{t}+d\sum_{j=1}^{M^{\alpha}_{t}}Z_{j}\,,\ \ X_{0}=x\in\mathbb{R}^{d}\,$, (26)
for a $d\times d$ matrix $\Sigma$, a $d$-dimensional Brownian motion $W$, a Poisson process $M^{\alpha}$ with intensity $\lambda^{\alpha}_{t}=\Lambda_{1}+\Lambda_{2}\|\alpha_{t}\|^{2}_{2}\,$ for $\Lambda_{1},\Lambda_{2}\geq 0$ and independent i.i.d. mean-zero square integrable $d$-dimensional random vectors $Z_{j}$ with $\mathbb{E}[\|Z_{j}\|^{2}_{2}]=\upsilon$, $j=1,2,\dots$ This problem admits a closed-form solution (see Appendix C), which provides a benchmark for our two numerical algorithms. We report their mean absolute errors with respect to the true solutions $V$ and $\alpha^{*}$, defined as ${\rm MAE}_{V}=\frac{1}{M}\sum_{i=1}^{M}|V_{\theta^{(k_{*})}}\hskip-1.05273pt(t_{i},x_{i})-V(t_{i},x_{i})|$ and ${\rm MAE}_{\alpha}=\frac{1}{M}\sum_{i=1}^{M}\|\alpha_{\phi^{(k_{*})}}\hskip-1.13809pt(t_{i},x_{i})-\alpha^{*}(t_{i},x_{i})\|_{2}$ on a test set of size $M$ uniformly sampled from $0,1]\times[-2.5,2.5]^{d}$.
@@
Figure 2(a) compares ${\rm MAE}_{V}$ and runtimes of GPI-PINN and GPI-CBU as functions of the number of epochs $k$ for 10-dimensional ($d=10$) LQR problems (25) with and without jumps. It can be seen that GPI-PINN and GPI-CBU exhibit similar convergence in the number of epochs. The residual-based approach of GPI-PINN makes it more stable than GPI-CBU; see also Baird (1995). On the other hand, GPI-CBU has a lower computational cost. This is already the case without jumps (left plot of Figure 2(a)) since it avoids the numerical evaluation of third-order derivatives but becomes much more significant with jumps (right plot of Figure 2(a)) as it also circumvents the numerical integration of the jumps.
@@
(a) Comparison of ${\rm MAE}_{V}$ (blue) and runtime in seconds (green) of GPI-PINN (solid line) and GPI-CBU (dashed line) for a 10-dimensional LQR problem without jumps (left) and with jumps (right).
@@
Figure 2 shows results of GPI-CBU applied to a 50-dimensional LQR problem with jumps compared to the analytical solution (30)-(33). While GPI-PINN is inefficient for this problem, GPI-CBU achieves high accuracy in the approximation of the value function as well as the optimal policy. Additional results for up to 150-dimensional LQR problems are reported in Appendix C.
@@
Figure 2: Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) at $t=0$ for $x=(x_{1},0,\dots,0)$ with $x_{1}\in[-2.5,2.5]$ for a 50-dimensional LQR problem with jumps. Orange dotted line: numerical results of GPI-CBU with $\pm 1$ standard deviation given by orange shaded area. Blue line: analytical solution (30)–(33).
@@
Figure 3 shows the accuracy of GPI-CBU on a 10-dimensional LQR problem with jumps compared with the two popular model-free RL algorithms PPO and SAC as well as the model-based discrete-time approach of Han & E (2016) applied to a time-discretization of the state dynamics (26). It can be seen that in this setup, PPO and SAC cannot compete with the two model-based approaches since they do not explicitly use the dynamics (26) but instead solely rely on sampling from the environment. The method of Han & E (2016) outperforms PPO and SAC but does not achieve the same accuracy as GPI-CBU due to discretization errors and since it does not generalize well to unseen points in the test set. Trying to cover the space-time domain well, we ran the method of Han & E (2016) from several randomly sampled starting points $x_{0}\in D$. But being a local method, it tends to learn the optimal control only along the optimal state trajectories $(t,X^{\alpha^{*}}_{t})_{0\leq t\leq T}$, which results in poor performance in parts of the space-time domain that are not explored well. Additional results are discussed in Appendix C.
@@ 5.2 Optimal consumption-investment with jumps
for two utility functions [^7] In our numerical experiments, we choose CRRA utility functions. $u,U\colon\mathbb{R}^{+}\to\mathbb{R}$, where $Y^{\alpha}_{t}$ is the wealth process evolving like
$\frac{dY^{\alpha}_{t}}{Y^{\alpha}_{t-}}=\left(r_{t}+\sum_{i=1}^{n}\pi^{i}_{t}(\mu^{i}-r_{t})-c_{t}\right)dt+\sum_{i=1}^{n}\pi^{i}_{t}\!\left[\sqrt{\sigma^{i}_{t}}\ \Sigma_{S,i}^{\top}\,dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\right]$
for a stochastic interest rate $r_{t}$, expected return rates $\mu^{i}$, stochastic volatilities $\sigma^{i}_{t}$ and stochastic jump intensities $\lambda^{i}_{t}$ whose dynamics are given in Appendix D.2 below. We consider strategies of the form $\pi^{i}(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$ and $c(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$. This problem has $d=2n+2$ state variables. The corresponding HJB equation, given in (42) in Appendix D.2, does not have an analytical solution, but Figure 4 shows that GPI-CBU converges numerically. More details about this consumption-investment problem with stochastic coefficients are provided in Appendix D.2.
In Appendix D.1, we also consider a simplified version of the problem where the coefficients $r_{t},\sigma^{i}_{t},\lambda^{i}_{t}$ are constant. This case can be reduced to a 1-dimensional ODE that can be solved with a standard Runge–Kutta scheme to obtain reference solutions. GPI-CBU generates results that are virtually indistinguishable from the Runge–Kutta results.
Figure 4: Losses $\widehat{\mathscr{L}}^{(k)}_{1}\hskip-1.42262pt(\theta^{(k+1)})$ (left) and $\widehat{\mathscr{L}}^{(k)}_{2}\hskip-1.42262pt(\phi^{(k+1)})$ (right) of GPI-CBU as a function of the epoch $k$. The blue curve in the left plot represents the interior loss of $\widehat{\mathscr{L}}^{(k)}_{1}$ and the orange curve its boundary part, see Eq. (21).
@@ 6 Conclusions, limitations and future work
In this paper, we have introduced two iterative deep learning algorithms for solving finite-horizon stochastic control problems with jumps. Both use an actor-critic architecture and train two neural networks to approximate the value function and optimal control, providing global solutions over the entire space-time domain without requiring simulation or discretization of the underlying state dynamics. Our first algorithm, GPI-PINN, works well for high-dimensional problems without jumps but becomes computationally infeasible in the presence of jumps. The second algorithm, GPI-CBU, leverages an efficient expectation-free iteration based on Proposition 4.1 which makes it particularly well-suited for high-dimensional problems with jumps. Both algorithms are model-based. As such, they outperform model-free RL methods in cases where the underlying state dynamics are known. The accuracy and scalability of the two algorithms has been demonstrated in different numerical examples.
@@ Appendix A Proofs
Letting $v(\cdot)=v\Big(t+\frac{h^{2}}{2n},x+\frac{h}{\sqrt{2}}\sigma_{i}(t,x,a)+\frac{h^{2}}{2n}\beta(t,x,a)\Big)$, the second-order derivative of $\psi(h):=\sum_{i=1}^{n}v(\cdot)$ is given by
$\psi^{\prime\prime}(h)=$ $\sum_{i=1}^{n}\Bigg(\partial^{2}_{tt}v(\cdot)\frac{h^{2}}{n^{2}}+\frac{h}{n}[\nabla_{tx}v(\cdot)]^{\top}\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)+\frac{1}{n}\partial_{t}v(\cdot)+\frac{1}{n}[\nabla_{x}v(\cdot)]^{\top}\beta(t,x,a)$
$\hskip 25.60747pt+\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)^{\hskip-1.99168pt\top}\hskip-2.84526pt\nabla^{2}_{x}v(\cdot)\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)$
$\hskip 25.60747pt+[\nabla_{tx}v(\cdot)]^{\top}\hskip-2.27621pt\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)\frac{h}{n}\Bigg)\,.$
@@
$\psi^{\prime\prime}(0)=\partial_{t}v(t,x)+\left[\nabla_{x}v(t,x)\right]^{\top}\hskip-2.84526pt\beta(t,x,a)+\frac{1}{2}\text{Tr}\left[\sigma\sigma^{\top}\hskip-1.42262pt(t,x,a)\nabla^{2}_{x}v(t,x)\right],$
@@
$\mathbb{E}[G_{\zeta}(t,Y_{t},Z_{1},V^{\alpha},\alpha(t,Y_{t}))\mid Y_{t}]=g(Y)$
@@
see e.g. Theorem 4.1.15 of Durrett (2019). This concludes the proof. ∎
@@ Appendix B DGM Architecture
Figure 5: DGM architecture for the value neural network with $L=3$ (i.e. 4 hidden layers).
Each DGM layer in Figure 5 for the value network $V_{\theta}$ is as follows
@@
where the number of hidden layers is $L+1$, and $\odot$ denotes element-wise multiplication. The DGM parameters for the value network are
@@
The number of units in each layer is $N$. $\sigma_{V}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is a twice-differentiable element-wise nonlinearity and $\sigma_{V_{out}}:\mathbb{R}^{N}\to\mathbb{V}$ is the output activation function, also twice-differentiable, where $\mathbb{V}\subseteq\mathbb{R}$ is chosen such as to satisfy possible restrictions on the value function’s output, inferred from the form of the stochastic control problem (e.g. non-negative value function). The control network is designed following the same DGM architecture with $\sigma_{\alpha_{out}}:\mathbb{R}^{N}\rightarrow A$. Throughout the numerical examples of the paper, we consider $L=3$, each one with $N\hskip-0.42677pt=\hskip-0.42677pt50$ neurons, see Figure 5. For the value network, we use $\tanh$ as the activation function $\sigma_{V}$ and softplus for $\sigma_{V_{out}}$. For the control network, we adopt $\tanh$ for $\sigma_{\alpha}$, while the output activation $\sigma_{\alpha_{\text{out}}}$ is linear in Example 5.1 and sigmoid in Example 5.2.
Unless otherwise stated, we take a number of sample points $M_{1},M_{2}$ equal to $256$, a number of gradient steps $N_{1},N_{2}$ equal to $64$, and a maximum number of epochs $k_{*}\leftarrow\min\{k_{*},1500\}$. In practice, we sample a new batch of size $M_{1}+M_{2}$ at each gradient step for Step 1 and 2 of Algorithms 1 and 2, covering a total batch size of $(N_{1}+N_{2})(M_{1}+M_{2})$ per epoch $k$. The network parameters are then updated using the Adaptive Moment Estimation with constant learning rates $\eta_{1},\eta_{2}$ equal to $0.001$. For both GPI-PINN and GPI-CBU, the algorithms are fairly robust to these choices. In our setting, the most critical hyperparameters for convergence are then the proportionality factors $\xi_{1},\xi_{2}$, which we determined using the approach of Wang et al. (2022), together with the scaling rate $\zeta$ (specific to GPI-CBU). In our experiments, we set $\zeta=1$ as it provides a good trade-off between convergence speed and accuracy of the improvements in GPI-CBU. Alternatively, one could consider an adaptive scaling factor $\zeta_{k}$ depending on the epoch $k$, similarly to learning rate schedules. On the other hand, we observe that negative scaling factors always fail to converge to the true solutions with exploding losses $\widehat{\mathscr{L}}_{1}^{(k)}$ and $\widehat{\mathscr{L}}_{2}^{(k)}$ after only a few epochs. Finally, Algorithms 1 and 2 are implemented using TensorFlow and Keras, which are software libraries in Python with reverse mode automatic differentiation. GPU acceleration has been used on a NVIDIA RTX 4090.

### v3

@@ 5 Numerical experiments
In our numerical experiments, we use the Deep Galerkin Method (DGM) architecture of Sirignano and Spiliopoulos (2018) for both the value and optimal control networks as it has been shown to empirically improve PINN performance. Details of the network design and hyperparameters are given in Appendix B.
@@ 5.1 Linear-quadratic regulator with jumps
$\inf_{\alpha}\mathbb{E}\left[\int_{0}^{T}c_{1}\|\alpha_{s}\|^{2}_{2}\,ds+c_{2}\|X^{\alpha}_{T}\|^{2}_{2}\right],\vskip-2.84526pt$ (18)
@@
where the infimum is over $d$-dimensional predictable processes $(\alpha_{t})_{0\leq t\leq T}$ and $X^{\alpha}$ is a $d$-dimensional process with controlled dynamics
$dX^{\alpha}_{t}=\alpha_{t}dt+BdW_{t}+d\sum_{j=1}^{M^{\alpha}_{t}}Z_{j}\,,\ \ X_{0}=x\in\mathbb{R}^{d}\vskip-2.84526pt$ (19)
for a $d\times d$ matrix $B$, a $d$-dimensional Brownian motion $W$, a Cox process $M^{\alpha}$ with intensity $\lambda^{\alpha}_{t}=\Lambda_{1}+\Lambda_{2}\|\alpha_{t}\|^{2}_{2}\,$ for constants $\Lambda_{1},\Lambda_{2}\geq 0$ and independent i.i.d. zero-mean square integrable $d$-dimensional random vectors $Z_{j}$ with $\mathbb{E}[\|Z_{j}\|^{2}_{2}]=\upsilon$, $j=1,2,\dots$
Problem (18) is of the general form (1) if instead of the infimum, one considers the supremum of the negative of the expectation. It can be reduced to a one-dimensional ODE which has a closed form solution in the special case $\Lambda_{2}=0$ and admits a very precise numerical solution via a standard ODE solver, such as Runge–Kutta, if $\Lambda_{2}>0$ (see Appendix C). This provides highly accurate reference solutions $V$ and $\alpha^{*}$ for the value function and optimal control.
@@
Figure 2(a) compares the performances of GPI-PINN and GPI-CBU on a 10-dimensional $(d=10)$ LQR problem of the form (18) with $T=1$, $B=I_{d}$, $c_{1}=1$, $c_{2}=1/4$ without jumps ($\Lambda_{1}=\Lambda_{2}=0$) and with jumps ($\Lambda_{1}=0$, $\Lambda_{2}=2$ and $d$-dimensional standard normal jumps $Z_{j}$). It shows mean absolute errors of $V_{\theta^{(k)}}$ with respect to $V$ given by ${\rm MAE}_{V}=\frac{1}{M}\sum_{i=1}^{M}|V_{\theta^{(k)}}\hskip-1.05273pt(t_{i},x_{i})-V(t_{i},x_{i})|$ on a test set of size $M$ uniformly sampled from $0,1]\times[-2.5,2.5]^{d}$ together with runtimes as functions of the number of epochs $k$. It can be seen that GPI-PINN and GPI-CBU exhibit similar convergence in the number of epochs. The residual-based approach of GPI-PINN makes it more stable than GPI-CBU; see also Baird (1995). On the other hand, GPI-CBU has a lower computational cost. This is already the case without jumps (left plot of Figure 2(a)) since it avoids the numerical evaluation of third-order derivatives but becomes much more significant with jumps (right plot of Figure 2(a)) as it also circumvents the numerical integration of the jumps.
@@
(a) Comparison of ${\rm MAE}_{V}$ (blue) and runtime in seconds (green) of GPI-PINN (solid lines) and GPI-CBU (dashed lines) for a 10-dimensional LQR problem (18) without jumps (left) and with jumps (right).
@@
Figure 2 shows results of GPI-CBU for the same LQR problem with jumps in $d=50$ dimensions. While the high dimensionality renders GPI-PINN infeasible, GPI-CBU achieves high accuracy in the approximation of the value function as well as the optimal policy. Additional results for up to 150-dimensional LQR problems are reported in Appendix C.
@@
Figure 2: Value function $V(0,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(0,x)$ (right) for $x=(x_{1},0,\dots,0)$ with $x_{1}\in[-2.5,2.5]$ for a 50-dimensional LQR problem with jumps. Orange dotted lines: numerical results of GPI-CBU with $\pm 1$ standard deviation given by orange shaded area. Blue lines: reference solutions $V$ and $\alpha^{*}$.
@@
Figure 3 shows the accuracy of GPI-CBU for the 10-dimensional LQR problem with jumps from the right panel of Figure 2(a) compared with the two popular model-free RL algorithms PPO and SAC as well as the model-based discrete-time approach of Han and E (2016) applied to a time-discretization of the state dynamics (19). It can be seen that in this setup, PPO and SAC cannot compete with the two model-based approaches since they do not explicitly use the dynamics (19) but instead only rely on sampling from the environment. The method of Han and E (2016) outperforms PPO and SAC but does not achieve the same accuracy as GPI-CBU due to discretization errors and since it does not generalize well to unseen points in the test set. Trying to cover the space-time domain well, we ran the method of Han and E (2016) from several randomly sampled starting points $x_{0}\in D$. But being a local method, it tends to learn the optimal control only along the optimal state trajectories $(t,X^{\alpha^{*}}_{t})_{0\leq t\leq T}$, which results in poor performance in parts of the space-time domain that are not explored well. Additional results are discussed in Appendix C.
@@ 5.2 Optimal consumption-investment with jumps
$\mathbb{E}\!\left[\int_{0}^{T}e^{-\rho s}u(c_{s}Y^{\alpha}_{s})ds+e^{-\rho T}U(Y^{\alpha}_{T})\right]$ (20)
for two utility functions [^7] In our numerical experiments, we choose CRRA utility functions. $u,U\colon\mathbb{R}_{+}\to\mathbb{R}$, where $Y^{\alpha}_{t}$ is the wealth process evolving like
$\frac{dY^{\alpha}_{t}}{Y^{\alpha}_{t-}}=\left(r_{t}+\sum_{i=1}^{n}\pi^{i}_{t}(\mu^{i}-r_{t})-c_{t}\right)dt+\sum_{i=1}^{n}\pi^{i}_{t}\!\left[\sigma^{i}_{t}\,dW^{i}_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\right]$ (21)
for a stochastic interest rate $r_{t}$, expected return rates $\mu^{i}$, stochastic volatilities $\sigma^{i}_{t}$, an $n$-dimensional Brownian motion $W$ with correlation matrix $\Sigma$, Cox processes $M^{i}$ with stochastic jump intensities $\lambda^{i}_{t}$ and random jump variables $Z^{i}_{j}$. We consider strategies of the form and $c(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$ and $\pi^{i}(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$. The problem has $d=2n+2$ state variables, consisting of $\sigma^{i}_{t},\lambda^{i}_{t},r_{t},Y^{\alpha}_{t}$ and $n+1$ decision variables $c_{t},\pi^{i}_{t}$, $i=1,\dots,n$. In general, the corresponding HJB equation, given in (35) in Appendix D.2, does not have an analytical solution. But it can be seen in Figure 4 that GPI-CBU converges numerically for $n=25$ financial assets. More details about this consumption-investment problem with stochastic coefficients are provided in Appendix D.2.
In Appendix D.1, we consider a simplified version of the problem where the coefficients $\sigma^{i}_{t},\lambda^{i}_{t},r_{t}$ are constant. This case can be reduced to a one-dimensional ODE that can be solved with a standard Runge–Kutta scheme to obtain reference solutions. GPI-CBU yields numerical solutions that are virtually indistinguishable from the Runge–Kutta results.
Figure 4: Training losses $\widehat{\mathscr{L}}^{(k)}_{1}\hskip-1.42262pt(\theta^{(k+1)})$ (left) and $\widehat{\mathscr{L}}^{(k+1)}_{2}\hskip-1.42262pt(\phi^{(k+1)})$ (right) of GPI-CBU as functions of the epoch $k$ for a consumption-investment problem of the form (20)–(21) with $n=25$ financial assets. The blue curve in the left plot represents the interior loss of $\widehat{\mathscr{L}}^{(k)}_{1}$ and the orange curve its boundary part; see (14).
@@ 6 Conclusions, limitations and future work
In this paper, we have introduced two iterative deep learning algorithms for solving finite-horizon stochastic control problems with jumps. Both use an actor-critic approach and train two neural networks to approximate the value function and optimal control, providing global solutions over the entire space-time domain without requiring simulation or discretization of the underlying state dynamics. Our first algorithm, GPI-PINN, works well for high-dimensional problems without jumps but becomes computationally infeasible in the presence of jumps. The second algorithm, GPI-CBU, leverages an efficient expectation-free updating rule based on Proposition 4.1 which makes it particularly well-suited for high-dimensional problems with jumps. Both algorithms are model-based. As such, they outperform model-free RL methods in cases where the underlying state dynamics are known. The accuracy and scalability of the two algorithms has been demonstrated in different numerical examples.
@@ Appendix A Proofs
Denoting $v_{i}(\cdot)=v\Big(t+\frac{h^{2}}{2n},x+\frac{h}{\sqrt{2}}\sigma_{i}(t,x,a)+\frac{h^{2}}{2n}\beta(t,x,a)\Big)$, the second-order derivative of $\psi(h):=\sum_{i=1}^{n}v_{i}(\cdot)$ is given by
$\psi^{\prime\prime}(h)=$ $\sum_{i=1}^{n}\Bigg(\frac{1}{n}\partial_{t}v_{i}(\cdot)+\frac{h^{2}}{n^{2}}\partial^{2}_{tt}v_{i}(\cdot)+2\frac{h}{n}\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)^{\top}\nabla_{tx}v_{i}(\cdot)$
$\qquad+\frac{1}{n}\beta^{\top}(t,x,a)\nabla_{x}v_{i}(t,x)+\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)^{\hskip-1.99168pt\top}\hskip-2.84526pt\nabla^{2}_{x}v_{i}(\cdot)\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)\Bigg).$
@@
$\psi^{\prime\prime}(0)=\partial_{t}v(t,x)+\beta^{\top}(t,x,a)\nabla_{x}v(t,x)+\frac{1}{2}\text{Tr}\left[\sigma\sigma^{\top}\hskip-1.42262pt(t,x,a)\nabla^{2}_{x}v(t,x)\right],$
@@
$\mathbb{E}[G_{\zeta}(t,Y_{t},Z_{1},V^{\alpha},\alpha(t,Y_{t}))\mid Y_{t}]=g(Y_{t})$
@@
see e.g. Theorem 4.1.15 of Durrett (2019). This completes the proof. ∎
@@ Appendix B DGM Architecture
Figure 5: DGM architecture of the value neural network with $L=3$ (i.e. 4 hidden layers).
Each DGM layer in Figure 5 in the value network $V_{\theta}$ is of the following form
@@
where the number of hidden layers is $L+1$, $\cdot$ denotes matrix multiplication and $\odot$ element-wise multiplication. The DGM parameters of the value network are
@@
The number of units in each layer is $N$. $\sigma_{V}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is a twice-differentiable element-wise nonlinearity and $\sigma_{V_{out}}:\mathbb{R}^{N}\to\mathbb{V}$ is the output activation function, also twice-differentiable, where $\mathbb{V}\subseteq\mathbb{R}$ is chosen so as to satisfy possible restrictions on the value function’s output, resulting from the form of the stochastic control problem (e.g. non-negative value function). The control network is designed following the same DGM architecture with $\sigma_{\alpha_{out}}:\mathbb{R}^{N}\rightarrow A$. Throughout the numerical examples of the paper, we use $L=3$ with $N=50$ neurons in each of the DGM layers, see Figure 5. In the value network, we use $\tanh$ for the activation function $\sigma_{V}$ and softplus for $\sigma_{V_{out}}$. For the control network, we adopt $\tanh$ for $\sigma_{\alpha}$, while as output activation $\sigma_{\alpha_{\text{out}}}$ we choose the identity in Example 5.1 and softplus in Example 5.2.
Unless stated otherwise, we use a number of sample points $M_{1},M_{2}$ equal to $256$, a number of gradient steps $N_{1},N_{2}$ equal to $64$ and a maximum number of epochs $k_{*}=1500$. The network parameters are updated using Adam (Kingma and Ba, 2014) with constant learning rates $\eta_{1},\eta_{2}$ equal to $0.001$. Both GPI-PINN and GPI-CBU are fairly robust with respect to these choices. In our setting, the most critical hyperparameters for convergence are the proportionality factors $\xi_{1},\xi_{2}$, which we determined using the approach of Wang et al. (2022), together with the scaling rate $\zeta$ (specific to GPI-CBU). In our experiments, we set $\zeta=1$ as it provides a good trade-off between convergence speed and accuracy of the improvements in GPI-CBU. Alternatively, one could consider an adaptive scaling factor $\zeta_{k}>0$ depending on the epoch $k$. On the other hand, negative scaling factors resulted in non-convergence in all our experiments with exploding losses $\widehat{\mathscr{L}}_{1}^{(k)}$ and $\widehat{\mathscr{L}}_{2}^{(k)}$ after only a few epochs.
Algorithms 1 and 2 were implemented using TensorFlow and Keras with GPU acceleration on an NVIDIA RTX 4090.